How do you solve this equation using the quadratic formula?
4(x-2)^2=20
Thanks :)

Found 2 solutions by dabanfield, Alan3354:
Answer by dabanfield(803)

How do you solve this equation using the quadratic formula?
4(x-2)^2=20

First expand the above and rewrite in the standard form ax^2 + bx + c = 0:
4(x-2)^2 = 20
4 = 20
4x^2 - 16x + 16 = 20
4x^2 - 16x - 4 = 0
Here we see a=4, b=-16 and c=-4. Now you can solve using the quadratic formula.

Answer by Alan3354(69443)

Try dividing both sides by 4.
Then what?